OpenStudy (anonymous):

You are ordering a pizza and soda for the big game. A deep dish pizza costs $10.25, and a two-liter of soda is $2.00. For each touchdown that the Eagles score on Sunday, Vinny's Pizza lowers the cost by $1.50. How many touchdowns will the Eagles have to score for your bill to come to $4.75? Explain how to solve for the missing value using an equation.

OpenStudy (nurali):

We can see that the price of the soda remains constant at 2.00. We want the bill to come to 4.75 thus we need the price of the pizza to be (4.75-2.00)=2.75. Now, the original pizza price was 10.25 so the price decrease is 10.25-2.75=7.5 For each touchdown, Vinny's lowers the cost by 1.50. So number of touchdowns required to get a price decrease of 7.5 is 7.5/1.50 = 5 touchdowns. Hope that helps.